OpenStudy (anonymous):

Write the following statement as an equation and solve. Jake Drewrey, owner of Jake's Landscape Service, knows that one of his fertilizer tanks holds twice as many gallons of liquid fertilizer as a second tank. The two tanks together hold 295 gal. If both tanks are filled to capacity, how many gallons of fertilizer does each tank hold?

OpenStudy (danjs):

Let A = Tank 1 Capacity B = Tank 2 Capacity one of his fertilizer tanks holds twice as many gallons of liquid fertilizer as a second tank. A = 2 * B two tanks together hold 295 gal A + B = 295

OpenStudy (danjs):

Solve for A and B, the capacity of each of the tanks in gallons of fertilizer. A = 2*B A+B = 295

OpenStudy (danjs):

Sub A = 2*B , into second equation 2B + B = 295 Solve for B (Gallons tank B can hold), then use that value in either of the equations, and solve for A ( Gallons tank A can hold).